OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

How do i start?

OpenStudy (anonymous):

i would multiply both sides of first eqn by -1 and then add to 2nd, which eliminates the x variable... so you are left with 9y -3z = 12 now multiply both sides of the 1st eqn by -2 and add to the 3rd, which eliminates x again, yielding -9y -z = 10 no with both these new eqns, add them which eliminates the y variable and you get -4z = 22 so z =-11/2

OpenStudy (anonymous):

*now with

OpenStudy (anonymous):

i dont understand...